

********************************************************************
UARS WEEKLY STATUS REPORT - 23 SEP 2005
********************************************************************
***********
UARS DE-BOOST ACTIVITY
***********
The first 18-minute thruster burn (first of 6 or 7 burns to
lower the UARS orbit), was originally planned for 29 September
2005. It is now being delayed until 4 October 2005.
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are waiting for DAAC archival.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. PEM data from
1991272 through 2005169 are available from SwRI. Users of SDDAS have this
data currently available. Please update your meta data to reflect the
current PEM data holdings.
Tests of the AXIS1 data file formats have been successful. The AXIS1
format design has been coded. A test file has been generated. A check
program with processing scripts have also been generated. Example binary
read program is now under development and will be tested on the generated
example binary data file.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG is participating in the on-going UARS end of mission
planning activities.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 September 2005 (Orbit 76625, GMT Day 253)
through 16 September 2005 (Orbit 76730, GMT Day 259).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 253/1926 76637 POWER OFF
The beta angle ranged between the angles of 32.1 degrees to 06.7
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 2 degrees on 17 September 2005 (DOY 260).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 4.20 to 7.32 degrees C. Additional
battery performance data for each battery on 09 September 2005 and 16
September 2005 are provided in Appendix D.
The clock error ranged from 12.0 to -10.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 13
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continues to closely monitor the condition of
battery 2 and battery 3. Battery 2 and 3 have not been sharing load
well since the cell failure on battery 2. Battery 3 carries roughly
3/4 of the power load while battery 2 provides the remaining 1/4. As
the beta angle goes lower, this discontinuity is slowly getting
exaggerated. NASA power branch has been providing daily support in
monitoring the UARS battery condition. The FOT has been authorized to
disable all platform and instrument heaters on the SSPP. This will be
done at the earliest available opportunity.
The FOT completed the planned 30-second burn on 14 September 2005.
The burn was started as planned at 257/1455 and terminated 30 seconds
later. Preliminary analysis revealed that the initial transients were
greater than expected. Therefore the FOT was not able to see any off
pulsing of the translation thrusters during the 30 seconds that the
thrusters were on for. It was decided to repeat this 30-second burn
again but this time with a BIAS in the pitch and yaw axis to induce
quicker response by the attitude thrusters and allow the FOT to see
the translation thrusters off pulse. 19 September 2005 was selected
as the date to execute this repeat burn. The burn is set to start
during the 257/1445 event with a fire time of 1455 UTC.
The FOT also updated TMON 45 and 46 with newer TMONs coded by our FSW
group. The updated TMONs also changed the way in which TMON 45
functions. TMON 45 use to check for battery 2 full charge counter to
be greater than a select time, and then take battery 2 off-line. This
worked well initially at the high angle that UARS was at, but it was
taking too long before battery 2 was being taken off relay and the C/D
ratio on battery 2 was higher than the power branch liked. TMON 45
now does the following: checks spacecraft time in day to be greater
than 5 minutes; checks to ensure battery 2 is on relay; checks to
ensure the state of charge (SOC) is greater than 99.5%. When these
conditions are met then battery 2 gets taken off relay. After a few
orbits of working with this new TMON, a change the SOC number was
made. The check is now for SOC greater than 99.8%. TMON 46 time in
night was increased to greater than 12.27 minutes as battery 3 was
charging battery 2 when it came back on relay.
MEETINGS
The FOT attended a De-boost status meeting this week on 12 September
2005. FDF predicted a SMA delta of ~590 meters for the 30 second
calibration burn that is slated for 14 September 2005.
The FOT attended a Command Authorization Meeting (CAM) on 13 September
2005. FDF presented the predicted orbital analysis for the 30-second burn.
The burn script that the FOT was to use was reviewed. There were no
showstoppers and the FOT was given permission to proceed with the
calibration burn on 14 September 2005.
The FOT attended the weekly Technical Working Group meeting on 15
September 2005. The less than expected performance of the 30-second
burn that took place on 14 September 2005 was discussed. It was
proposed to attempt a repeat of the 30-second burn on the following 19
September 2005, but this time with a BIAS of -0.75 degrees in pitch
and 1.55 degrees in yaw.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 September 2005 (DOY 260).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Another 4-thruster calibration burn of 30 seconds in length is
scheduled for 19 September 2005 (DOY 262).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 September 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
76637 253/1926 05-318 HALOE POWER OFF
76665 255/1455 05-319 OBC TMON 45 UPDATE
76665 255/1550 05-320 OBC OBC FSW TO 69A
76666 255/1634 05-321 OBC TMON 45 UPDATE
76672 256/0223 05-322 MPS VT L-5 TO VT L-6
76681 256/1630 05-323 OBC TMON45/46 - CHANGE TIMING (10:00)
76694 257/1255 05-324 MACS ENGINEERING BURN (STEP 1)
76694 257/1411 05-324 MACS ENGINEERING BURN (STEP 3,5,6)
76695 257/1447 05-324 MACS ENGINEERING BURN (STEP 7-24)
76695 257/1542 05-325 OBC MEMMON, DUMPS
76696 257/1627 05-326 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76697 257/1806 05-327 OBC TMON 46 TIMING CHANGE (12:20)
76697 257/1904 05-328 OBC OPEN EPH THRESHOLD
76699 257/2051 05-329 OBC OPEN EPH THRESHOLD/RELOAD
76700 257/2246 05-329 OBC SET DEFAULT EPH THRESHOLD
76715 258/2336 05-330 OBC ATC NOOP & XMTR OFF
76718 259/0422 05-331 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 September 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 September 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
76696 257/1635 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
76718 259/0427 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 September 2005
09 September 2005 - GMT Day 252 Beta = 34.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.74/ -3.68 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.87/ +2.91 +24.0/-17.6 24.5 1.05 11.5
3 179.2/-291.2 +6.43/ +5.14 +29.6/-23.2 24.5 1.01 24.3
16 September 2005 - GMT Day 259 Beta = 06.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/ -3.68 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.87/ -0.23 +21.6/-13.6 24.3 1.03 7.1
3 067.2/-033.6 +8.07/ +7.09 +49.6/-22.8 24.2 1.01 28.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
76672 256/0223 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 September 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-044 FOT See List Sept 05 30832 Generic Late Acquisition